A 10 OUNCES SOLUTION OF ALCOHOL AND IODINE IS 3% IODINE.HOW MUST ALCOHOL MUST BE ADDED TO MAKE A SOLUTION FOR TREATING CUTS THAT IS ONLY 2% IODINE?
GOOD MORNING SIR/MA'AM.I NEED A STEP-BY-STEP SOLUTION. THANK YOU


Answer by mananth(16946)

--------percent---------------- quantity
Iodine 3 ---------------- 10 ounces
alcohol 0 ---------------- x ounces
Total 2 ---------------- 10 + x ounces

3*10+0*x =2(10+x)
30+0x=20 +2x

-2x=20-30

-2x =-10

/-2

x=5 ounces alcohol
m.ananth@hotmail.ca